OpenStudy (anonymous):

trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 mph and train B at 44 mph. Train A passes a station at 1:2-am. If train B passes the same station at 1:50am, at what time will train B catch up to train A?

OpenStudy (anonymous):

that's supposed to be 1:20am not 1:2-am

OpenStudy (anonymous):

Using the line format: \[y=mx+b\] You solve for the two equations: \[y_{a} = 0, x_{a} = 1.333\]\[y_{a} = 40x_{a} + b\]\[0 = 40(1.333) + b\]\[-53.333 = b\]\[y_{a} = 40x - 53.333\] \[y_{b} = 0, x_{b} = 1.8333\]\[y_{b} = 44x_{b} + b\]\[0= 44(1.8333) + b\]\[-73.333 = b\]\[y_{b} = 44x -73.333\]You then solve the two equations: \[40x - 53.333 = 44x - 73.333\]\[20 = 4x\]\[x = 5\]If I didn't mess up horribly somewhere, that means they should meet at 5:00 am.

OpenStudy (anonymous):

You can calculate the distance train A has moved since it passed the station untill train B passes that station, that will be 20 miles, as the time diffenrence is a half hour. If we say the train B catches up with 4 mph, then divide the distance with the speed, 20miles / 4 mph = 5 hours, 5 hours plus 1:50 am = 6:50 am

OpenStudy (anonymous):

r t = d Solve 40 (t + 1/2 ) = 44 t, for t t = 5 hours time of day = 1:50AM +5 hours or 6:50AM